The density of a gas is 1.964 g dm-3 at 273 K and 76 cm Hg. The gas is

Grade:11

1 Answers

Arun
5 years ago
Dear student
 
Density=PM/RT==> M=dRT/P==>M=(1.964×10^(-3) × 82.1×273)/1==>M=44gTherefore, the gas is co2.76cmHg is approximately equal to 1 bar. Therefore, pressure taken is 1 bar.